You would like to estimate the weighted average cost of capital for a new airline business. Based on its industry asset​ beta, you have already estimated an unlevered cost of capital for the firm of 9 % . ​However, the new business will be 26 % debt​ financed, and you anticipate its debt cost of capital will be 7 % . If its corporate tax rate is 38 % ​, what is your estimate of its​ WACC?

The equity cost of capital is %. (Round to one decimal​ place.)
Answer #1

WACC = we*Re + wd*Rd*(1-Tc)

where,

wd = weight the firm’s debt = 0.26
we = weight firm’s equity = 1 - 0.26 = 0.74
Re = cost of equity = 0.09
Rd = cost of debt = 0.07
Tc = corporate tax rate = 0.38

=> WACC = 0.74*0.09 + 0.26*0.07*(1-0.38) = 0.0779 or 7.79%